Find fx(x,y), fy(x,y), fx(1,3), and fy(-2,4) for the given function. If 

𝑧 = 𝑓(𝑥, 𝑦) = 3𝑥ଷ𝑦ଶ − 𝑥ଶ𝑦ଷ + 4𝑥 + 9



1
Expert's answer
2021-06-12T00:47:26-0400

Given f(x,y) = 3x2y - xy2 + 4x + 9

Now fx(x,y) = (d/dx)(3x2y - xy2 + 4x + 9)

= 6xy - y2 + 4

Now fx(1,3) = 6*1*3 - 32 + 4 = 13

Now fy(x,y) = (d/dy)(3x2y - xy2 + 4x + 9)

= 3x2 - 2xy

Now fy(-2,4) = 3(-2)2 - 2(-2)(4) = 28